Five balls are selected at random without replacement from an urn containing three white balls and five blue balls. Find the probability that two or three of the balls are white. (Round your answer to three decimal places.)
Answer by scott8148(6628)

number of ways to select 5 from 8 ___ 8C5 = 56

number of ways to select 2 white from 3 ___ 3C2 = 3

number of ways to select 3 white from 3 ___ 3C3 = 1

probability of 2 or 3 white ___ (3 + 1) / 56 = 4 / 56 = 1 / 14